Find 1/2 of 1%190 tons
225 is 25% of what amount?
120 is what percent of 240?
what do discounts of 25%, 8%, and 5% amount to on a bill of $300.00?

Answer by melanie125(2)

225 is 25% of what amount? n= 225/25%= 900
120 is what percent of 240? n= 120/240 x 100= 50%